How do you go about showing work and solving this problem o = y -11 - 2y
Answer by unlockmath(1688)

Hello,
First we want to get y all by itself. To do this, combine like terms and add 11 to both sides to show like this:
o+11 = -11 +11 - y (Notice 11 cancels on the right leaving -y)
11=-y (Now we need to get rid of the negative sign in front of the y so we multiply both sides by -1.
11(-1)=-1(-y) This can be rewritten as:
-11=y
Check it out by plugging -11 into the equation. You're good to go.
RJ Toftness
www.math-unlock.com